# The ML Supply Chain in the Era of Software 2.0: Lessons Learned from Hugging Face

### v1

@@ 2. Background and Related Work
In conclusion, while our study has been informed by prior work (Castaño et al., 2024a; Yang et al., 2024; Jiang et al., 2024a, 2023d; Oreamuno et al., 2024; Jiang et al., 2023c; Pepe et al., 2024; Taraghi et al., 2024; Jiang et al., 2023b), our investigation focuses on the structure and characteristics of the ML supply chain itself and the challenges encountered in managing it. More specifically, while previous work examined deficiencies in documentation and licensing at the level of individual models, we investigate the ML supply chain at a higher level by examining how it is structured and how models relate to each other.
@@ 1. RQ0:
    What documentation deficiencies exist in Hugging Face model and data cards that potentially complicate mapping the ML supply chain? To understand the ML supply chain, we needed to collect and analyze Hugging Face metadata. Using this data, we identified several issues and challenges related to documentation that introduced hurdles to obtaining a comprehensive understanding of the ecosystem. We examine these challenges in detail because with poorly and inconsistently described models and datasets, developers may struggle to leverage the ML supply chain effectively. This preliminary research question discusses encountered examples of noisiness issues occurring in Hugging Face model and dataset documentation.
@@ 2. RQ1:
    What is the structure of the ML supply chain? The ML supply chain, like any other supply chain, is characterized by dependencies across components. This research question looks into the complexities of such a dependency graph.
@@ 3. RQ2:
    What is the current licensing landscape for models and datasets and what potential compliance challenges does it pose? ML models are software, so software licenses govern their use and redistribution. As such, these models present new challenges in the area of license compliance. To understand these challenges, we examine the state of licensing in the ML supply chain by 1) analyzing common licenses and license categories, 2) exploring licensing decisions made in the supply chain, and 3) investigating potential compliance challenges, such as incompatibilities between parent/child model licenses, ambiguities introduced by undeclared, unknown, or custom licenses, and uncertainty surrounding how license terms apply in context.
@@
In the following subsections, we describe the study methodology, which is depicted in Fig. 1. We start by mining model and dataset lists. Next, for each model/dataset, we extract information necessary to construct the ML supply chain graph, recording and analyzing documentation deficiencies encountered during the process (RQ0). Then, we systematically analyze and evaluate the structure of the resulting supply chain graph (RQ1). This same graph, along with licensing information, is then used to address RQ2.
@@ 3.1. Data collection
To answer the research questions, we used the Hugging Face API to extract a list of all models publicly hosted on the platform. An initial list of 500,000 models was pulled on July 9, 2024, and a full list of all available models was obtained on July 11, 2024. Using the same APIs, we mined data (model cards and metadata) for each model on this list. For models that were unavailable through the API (as described in Section 4.1), we employed a web-scraping technique relying on a combination of Python’s request module and the BeautifulSoup HTML parser. Mining took place between July 9 and July 12, 2024, yielding 760,460 mined models. We also used the Hugging Face API to gather the list of all datasets publicly hosted on the platform and available through the API as of July 9, 2024. We make all the raw data acquired from mining available in our replication package (Stalnaker et al., 2025).
While a handful of research datasets for Hugging Face models already exist (Jiang et al., 2024b, 2023d), we opted to create our own dataset for this work for three reasons. First, we aimed to analyze the most current data, which includes new models and model families (e.g., new models recently added to the Llama family of models) and changes in licensing frequencies (see Section 6). Second, we aimed to collect and analyze the most comprehensive and complete dataset; as shown in Table 1, our dataset is significantly larger than existing datasets. Third, unlike Jiang et al. (Jiang et al., 2024b) and Pepe et al. (Pepe et al., 2024), we were not focused on model-to-GitHub repository dependencies; accordingly, we aimed to design a dataset tailored for mapping model-to-model dependencies.
Table 1. Comparison of Hugging Face (HF) datasets used in prior work
@@
(Yang et al., 2024) - 24,065 Explore the state of dataset documentation on HF
(Jiang et al., 2024a) 500 - Investigate PTM* naming practices/anomolies on HF
@@
(Jiang et al., 2023b) / (Jiang et al., 2024b) 281,638 - Create a dataset examining relationship between PTMs and OSS
@@ 3.2. Data normalization and cleaning
Once the data for models and datasets was downloaded, it needed to be cleaned, standardized, and stored in a more usable form. This process involved standardizing the names of declared base models and datasets and mapping to unique identifiers where possible. For example, xlm-roberta-base is a shorthand for FacebookAI/xlm-roberta-base, which has the unique internal ID 621ffdc036468d709f174364. We relied on Hugging Face’s internal mechanisms to resolve these differences where possible. For example, attempting to load the model page for xlm-roberta-base on Hugging Face redirects to the page for FacebookAI/xlm-roberta-base. Additionally, we resolved inconsistencies in field inputs. For example, [], “”, and None were all being used to denote no declared licenses. We standardized these variations to the empty list to make later computation simpler. We further elaborate on documentation challenges that required special attention and resolution in our discussion of RQ0.
@@ 3.3. Extracting licensing information
We extracted licensing information from the tags present in model metadata. While licenses can also be declared in the CardData attribute of the model metadata, we found through an analysis of our dataset that the tags were more complete in every instance than the CardData attribute. In nearly all instances, licensing information was in both locations (99.9%). However, for 0.1% of models, the licensing information was exclusively in the tags, and there were no instances that relied only on the CardData attribute. We also looked for discrepancies in those cases where both locations were utilized. We identified only 134 such discrepancies, and in all cases, an additional license was declared in the tags while it was not declared in the CardData. There were no irreconcilable differences (i.e., MIT vs GPL). From this, we concluded that the model tags are the most complete machine-readable source for licensing information.
Although licensing information can also be found in other, harder to programmatically extract sections of a model card, we sought to further confirm that reliance on tags to identify the declared license used by the model was justified. Accordingly, one author manually reviewed the model cards for the top 100 models (sorted by number of downloads) marked as having no license in the tags. We anticipated that these 100 models were more likely to have complete and reliable documentation because of their popularity as compared with other models in our dataset, consistent with the findings of Jones et al. (Jones et al., 2024). From this sample of 100 models, we identified only 16 models that had declared licensing information elsewhere. Of those, one had licensing information added to its tags after our initial mining, bringing the total to 15 models. For nine of these models, the license could only be found transitively by following a link to the model’s associated GitHub page. Table 2 provides a more detailed breakdown of where the licensing information was located for each example.
@@
Detecting each model’s declared license would involve more complicated mining techniques, such as those that could traverse external links, and would require NLP-based approaches that extract licensing information from the text. These NLP approaches would have to not only recognize license names, which may not be in a standardized form but also pattern-match embedded license text. An LLM tool, like the one employed by Jiang et al. (Jiang et al., 2024b), could potentially be used to extract licensing information, but at a high cost (between \$0.01 and \$0.03 per model— or between \$7,000 and \$22,800 for our dataset). To ensure that our analysis captured the most accurate information possible and to better understand the formal documentation structure of Hugging Face, rather than employ such an approach, we focus on explicitly declared licenses. Given the likelihood that licenses are declared in tags, the infeasibility of manual annotation for all examples, and the inherent difficulties of automatically obtaining license information from anywhere in a model card, our reliance on tags for license detection appears to be an acceptable approach. Additionally, existing supply chain analysis tools rely heavily on machine-readable metadata, so we further justify that this is the appropriate locus of analysis given that the ultimate audience for such tools is the end user.
@@ 3.4. Data analysis
Once the dataset had been cleaned, to address RQ1, we used the networkx Python library (networkx, 2024) to create a directed graph representing the observable ML supply chain. Each node in the graph represents a distinct model, and an edge represents a dependency relationship. For example, if model $A$ listed model $B$ as a base model, there would be an edge from $B$ to $A$. We would describe model $B$ as the parent of model $A$ and model $A$ as the child of model $B$. Due to the inherent issues and incompleteness of model documentation, this graph, by its nature, only represents a subset of the total ML supply chain on Hugging Face.
For the sake of analysis, we flatten the individual model supply chains by extracting all longest paths (e.g., lineage chains) from the graph that go from source nodes to sinks. Each of these paths goes from a root base model to a model at the end of the ML supply chain (i.e., a leaf node with no dependents or outgoing connections).
Using the paths resulting from RQ1, we addressed RQ2 by analyzing the state of licensing in the ML supply chain, developing scripts and heuristics to look for parent/child license differences, and potential license compliance problems related to model/dataset dependencies. We define a parent/child license difference to refer to the difference between the license(s) of a parent and a child model/dataset. Assume, for example, that model $B$ is the child of model $A$. If model $B$ was licensed under MIT and model $A$ under Apache-2.0, we would say there is a parent/child license difference between $A$ and $B$. We avoid stronger terms such as “inconsistencies” and “incompatibilities”, since some of these differences may in fact be consistent and compliant with licensing terms. Two authors reviewed all distinct licenses in our dataset and grouped them into six categories, which can provide the basis for future research into license compatibility and licensing trends in the ML supply chain.
@@ 4. RQ0: Documentation challenges encountered on Hugging Face
In this section, we outline and discuss documentation issues that were observed on the Hugging Face platform while mapping connections in the ML supply chain. Specifically, this preliminary RQ outlines the observed shortcomings and challenges to 1) elaborate potential pitfalls future researchers may face in mining Hugging Face and 2) to more thoroughly report documentation issues extant on Hugging Face that have been left unexplored by prior work.
@@ 4.1. Inability to Access Metadata
Not all model metadata was accessible through the Hugging Face API. There were 7,258 model cards (0.95%) that had to be scraped using a combination of Python’s requests (requests, 2024) and BeautifulSoup (beautifulsoup, 2025) libraries. The most frequent challenge (i.e., found in 99.4% of the cases) resulted from gating based on the acceptance of terms and conditions (see Figure 2). This was particularly the case for models owned by or relying on models created by larger entities such as Google or Meta. Twenty-two models and their metadata were gated behind age restrictions imposed by the model owner [^1] Throughout the paper, by ”model owner”, we refer to the entity that uploaded the model to Hugging Face. We make no claims as to whether that entity had the legal right to do so or if they are the original author/creator of the model., requiring a Hugging Face account to access. Accessing the full repository, and thus metadata, required logging in with a Hugging Face account and then, in most cases, providing an email address. This hurdle to even access the metadata for a model could potentially complicate the creation of fully automated analyzers and ML/AIBOM (Machine Learning/Artificial Intelligence Bill of Materials) generators for an ML supply chain.
@@ 4.2. Incomplete Metadata
We observed that the documentation provided by model owners on Hugging Face is often incomplete. This is a trend not just for smaller, infrequently used projects but also for large projects with tens of thousands of downloads. For example, OpenAI’s clip-vit-large-patch14-336 model, with 5,827,027 downloads, has a model card that is almost entirely incomplete, save for basic hyper-parameter and framework version information (OpenAI, [n. d.]). Only 37.8% of models and 27.6% of datasets declared any kind of licensing information in a machine-readable way. We discuss the prevalence of declared components further in Section 5.
Some models may provide information on training data and architecture by linking to scientific papers that include these details, but such a practice introduces an additional hurdle in easily and programmatically obtaining that data. Ultimately, the reason why many models do not even mention the datasets used would require further investigation, including interviewing/surveying developers, which is out of the scope of this work.
@@
4.3. Additional Metadata Issues
4.3.1. The ”unknown” license.
The Hugging Face platform’s documentation provides a list of recognized licenses each with a unique, standardized short-hand identifier that can be utilized by users when creating their model/data cards (Face, 2024a). This list includes common Open Source Software (OSS), Creative Commons (CC), and ML-specific licenses as well as an “Other” catch-all category which encompasses less common, custom, or modified licenses not included in Hugging Face’s predefined license list An “unknown” licensing option is also provided, but there is no guidance provided as to when this option should be selected. Future work will need to explore when and why model owners select this option, particularly given that acknowledging that the license is “unknown” suggests, at best, a lack of due diligence or understanding on the part of the owner or, at worst, that the model/dataset raises copyright infringement issues. We observed 4,419 (1.5%) models and 2,194 (4.5%) datasets that used this “unknown” license tag, which inherently raises compliance challenges for dependent models.
4.3.2. Naming problems.

### v2

@@ 3.4. Summary
In conclusion, while our study builds on the basis of prior work (Castaño et al., 2024a; Yang et al., 2024a; Jiang et al., 2024b, 2023d; Oreamuno et al., 2024; Jiang et al., 2023c; Pepe et al., 2024; Taraghi et al., 2024; Jiang et al., 2023b), our investigation focuses on the structure and characteristics of the ML supply chain itself and the challenges encountered in managing it. More specifically, while previous work examined deficiencies in documentation and licensing at the level of individual models, we investigate the ML supply chain at a higher level by examining how it is structured and how models relate. Additionally, where previous work has examined the relationships between projects and models, we examine the relationships among multiple models. Thus, our work is beneficial not only to software developers, as is the case for prior work, but also to data scientists, ML engineers, and anyone who wants to reuse models with complex dependencies. By providing an ecosystem-level examination of the ML supply chain, we provide these groups with the information necessary to understand how models relate to one another and evolve, licensing concerns regarding ML components on Hugging Face, and potential challenges one can face regarding model documentation.
@@ 1. RQ0:
    What documentation deficiencies exist in Hugging Face model and data cards that potentially complicate mapping the ML supply chain? To understand the ML supply chain on Hugging Face, we needed to collect and analyze Hugging Face metadata. Using these data, we identified several issues and challenges related to documentation that introduced hurdles to obtaining a comprehensive understanding of the ecosystem. We examine these challenges in detail because with poorly and inconsistently described models and datasets, developers may struggle to leverage the ML supply chain effectively. This preliminary research question discusses examples of noisiness issues occurring in the Hugging Face model and the data set documentation.
@@ 2. RQ1:
    What is the structure of the ML supply chain? The ML supply chain, like any other supply chain, is characterized by dependencies across components. This research question looks at the complexities of such a dependency graph.
@@ 3. RQ2:
    What is the current licensing landscape for models and datasets and what potential compliance challenges does it pose? ML models are software, so software licenses govern their use and redistribution. As such, these models present new challenges in the area of license compliance. To understand these challenges, we examine the state of licensing in the ML supply chain on Hugging Face by 1) analyzing common licenses and license categories, 2) exploring licensing decisions made in the supply chain, and 3) investigating potential compliance challenges, such as incompatibilities between parent/child model licenses, ambiguities introduced by undeclared, unknown, or custom licenses, and uncertainty surrounding how license terms apply in context.
@@
In the following subsections, we describe the study methodology, which is depicted in Fig. 1. We start by mining model and dataset lists. Next, for each model/dataset, we extract information necessary to construct the ML supply chain graph, recording and analyzing documentation deficiencies encountered during the process ($RQ_{0}$). Then we systematically analyze and evaluate the structure of the resulting supply chain graph ($RQ_{1}$). The same graph, along with the licensing information, is then used to address $RQ_{2}$.
@@ 4.1. Data collection
To answer the research questions, we used the Hugging Face API to extract a list of all models publicly hosted on the platform. An initial list of 500,000 models was pulled on July 9, 2024, and a full list of all available models was obtained on July 11, 2024. Using the same APIs, we mined data (model cards and metadata) for each model on this list. For models that were unavailable through the API (as described in Section 5.1), we employed a web-scraping technique relying on a combination of Python’s request module and the BeautifulSoup HTML parser. Mining took place between July 9 and July 12, 2024, yielding 760,460 mined models. We also used the Hugging Face API to gather the list of all datasets publicly hosted on the platform and available through the API as of July 9, 2024. For each model in our dataset, information on base models and datasets was obtained, when available, from the CardData attribute of the model metadata as well as from the model tags. We make all the raw data acquired from mining available in our replication package (Stalnaker et al., 2025).
Although there are a handful of research datasets for Hugging Face models and datasets already exist (Jiang et al., 2024c, 2023d), we opted to create our own dataset for this work for three reasons:
  1. (1)
    We aimed to analyze the most current data, which includes new models and model families (e.g., new models recently added to the Llama family of models) and changes in licensing frequencies (see Section 7). The presence of these new models is significant given that many of these were the first open-weight models capable of competing with the, until then, assumed more powerful and untouchable proprietary options (e.g., OpenAI’s Chat-GPT, Google’s Gemini, and Anthropic’s Claude). The practical utility of these new models represented a paradigm shift for fine-tuning, reusing, and remixing existing open-weight models. Their inclusion is essential to understand the present state of the ML supply chain on Hugging Face.
  2. (2)
    We aimed to collect and analyze a more comprehensive and complete dataset; as shown in Table 1, our dataset is significantly larger than most existing datasets and comprises both datasets and models.
  3. (3)
    Unlike Jiang et al. (Jiang et al., 2024c) and Pepe et al. (Pepe et al., 2024), we were not focused on model-to-GitHub repository dependencies; accordingly, we aimed to design a dataset tailored for mapping model-to-model and dataset-to-model dependencies.
Table 1. Comparison of study goals and Hugging Face (HF) datasets used in prior work
@@
(Yang et al., 2024a) - 24,065 Explore the state of dataset documentation on HF
(Jiang et al., 2024b) 500 - Investigate PTM* naming practices/anomolies on HF
@@
(Jiang et al., 2023b) / (Jiang et al., 2024c) 281,638 - Create a dataset examining relationship between PTMs and OSS
@@ 4.2. Data normalization and cleaning
Once the data for models and datasets were downloaded, the data needed to be cleaned, standardized, and stored in a more usable form. This process involved standardizing the names of declared base models and datasets and mapping to unique identifiers where possible. For example, xlm-roberta-base is a shorthand for FacebookAI/xlm-roberta-base, which has the unique internal ID 621ffdc036468d709f174364. We relied on Hugging Face’s internal mechanisms to resolve short names where possible. For example, attempting to load the model page for xlm-roberta-base on Hugging Face redirects to the page for FacebookAI/xlm-roberta-base. Additionally, we resolved inconsistencies in field inputs. For example, [], “”, and None were all being used to denote no declared licenses. We standardized these variations to the empty list to make later computations simpler. We further elaborate on documentation challenges that required special attention and resolution in our discussion of RQ0.
@@ 4.3. Extracting licensing information
We extracted licensing information from the tags present in model metadata. While licenses can also be declared in the CardData attribute of the model metadata, we found through an analysis of our dataset that the tags were more complete in every instance than the CardData attribute. In nearly all instances where licensing information was present, it was found in both locations (99.9%). However, for 0.1% of models, the licensing information could be found exclusively in the tags, and there were no instances that relied only on the CardData attribute. We also looked for discrepancies in the cases where both locations were utilized. We identified only 134 discrepancies, and in all cases, an additional license was declared in the tags that was not declared in the CardData. As there were no irreconcilable differences (i.e., the tags declared an MIT license while the CardData attribute declared a GPL license), we concluded that the model tags are the most complete machine-readable source for licensing information.
Although licensing information can also be found in sections of a model card that are more difficult to programmatically extract from, we sought to further confirm that reliance on tags to identify the declared license used by the model was justified. Accordingly, one author manually reviewed the model cards for the top 100 models (sorted by number of downloads) marked as having no license in the tags. We anticipated that these 100 models were more likely to have complete and reliable documentation because of their popularity as compared with other models in our dataset, consistent with the findings of Jones et al. (Jones et al., 2024). From this sample of 100 models, we identified only 16 models that had declared licensing information elsewhere. Of those, one had licensing information added to its tags after our initial mining, bringing the total to 15 models. For nine of these models, the license could only be found transitively by following a link to the model’s associated GitHub page. Table 2 provides a more detailed breakdown of where the licensing information was located for each example.
@@
Detecting license declarations present in other sources (e.g., plain text or links to external sites) would involve more complicated mining techniques, such as those that could traverse external links, and would require NLP-based approaches that extract licensing information from the text. These NLP approaches would have to not only recognize license names, which may not be in a standardized form, but also pattern-match embedded license text. An LLM tool, like the one employed by Jiang et al. (Jiang et al., 2024c), could potentially be used to extract licensing information, but at a high cost (between \$0.01 and \$0.03 per model—or between \$7,000 and \$22,800 for our dataset). Relatedly, relying on a cheaper but smaller, locally hosted LLM would risk introducing hallucinated license information and so still require laborious manual human oversight. To ensure that our analysis captured the most accurate information possible and to better understand the formal documentation structure of Hugging Face, rather than employ such an approach, we focus on explicitly declared licenses. Given the likelihood that licenses are declared in tags, the infeasibility of manual annotation for all examples, and the inherent difficulties of automatically obtaining license information from anywhere in a model card, our reliance on tags for license detection appears to be an acceptable approach. Additionally, existing supply chain analysis tools rely heavily on machine-readable metadata, so we further justify that this is the appropriate locus of analysis given that the ultimate audience for such tools is the end user.
@@ 4.4. Data analysis
Once the dataset had been cleaned, to address RQ1, we used the networkx Python library (networkx, 2024) to create a directed graph representing the observable ML supply chain. Due to the inherent issues and incompleteness of model documentation, this graph, by its nature, only represents a subset of the total ML supply chain on Hugging Face. Each node in the graph represents a distinct model, and an edge represents a dependency relationship. For example, if model $A$ listed model $B$ as a base model, there would be an edge from $B$ to $A$. We would describe model $B$ as the parent of model $A$ and model $A$ as the child of model $B$. That is, model $A$ is derived from model $B$.
For the sake of analysis, we flatten the individual model supply chains by extracting all longest paths (i.e., lineage chains) from the graph that go from source nodes to sinks. Each of these paths goes from a root base model to a model at the end of the ML supply chain (i.e., a leaf node with no dependents or outgoing connections— no other models are derived from this model).
Using the paths resulting from RQ1, we addressed RQ2 by analyzing the state of licensing in the ML supply chain, developing scripts and heuristics to look for parent/child license differences and potential license compliance problems related to model/dataset dependencies. We define a parent/child license difference to refer to the difference between the license(s) of a parent and a child model/dataset. Assume, for example, that model $B$ is the child of model $A$. If model $B$ was licensed under MIT and model $A$ under Apache-2.0, we would say there is a parent/child license difference between $A$ and $B$. We avoid stronger terms such as “inconsistencies” and “incompatibilities,” since some of these differences may in fact be compatible with licensing terms. This is consistent with prior work (Duan et al., 2024) that found that the application of licensing terms, particularly for existing licenses, in the novel ML context can introduce uncertainties and ambiguities in terms of legal interpretation. Two authors reviewed all distinct licenses in our dataset and grouped them into six categories, which can provide the basis for future research into license compatibility and licensing trends in the ML supply chain.
@@ 5. RQ0: Documentation challenges encountered on Hugging Face
In this section, we outline and discuss documentation issues that were observed on the Hugging Face platform while mapping connections in the observed ML supply chain. Specifically, this preliminary RQ outlines the observed shortcomings and challenges to 1) elaborate potential pitfalls future researchers may face in mining Hugging Face and 2) to more thoroughly report documentation issues extant on Hugging Face that have been left unexplored by prior work.
@@ 5.1. Inability to Access Metadata
Not all model metadata was accessible through the Hugging Face API. There were 7,258 model cards (0.95%) that had to be scraped using a combination of Python’s requests (requests, 2024) and BeautifulSoup (beautifulsoup, 2025) libraries. The most frequent challenge, found in 99.4% of these cases, resulted from gating based on the acceptance of terms and conditions (see Figure 2). This was particularly the case for models owned by or relying on models created by larger entities such as Google or Meta. Twenty-two models and their metadata were gated behind age restrictions imposed by the model owner [^1] Throughout the paper, by ”model owner”, we refer to the entity that uploaded the model to Hugging Face. We make no claims as to whether that entity had the legal right to do so or if they are the original author/creator of the model., requiring a Hugging Face account to access. Accessing the full repository, and thus metadata, required logging in with a Hugging Face account and then, in most cases, providing an email address. These hurdles to accessing the metadata for a model could potentially complicate the creation of fully automated analyzers and ML/AIBOM (Machine Learning/Artificial Intelligence Bill of Materials) (Stalnaker et al., 2024; Xia et al., 2023a; Barclay et al., 2019, 2022) generators for an ML supply chain, which we believe should not rely on scraping potentially changing HTML content or on Selenium-based web-drivers for form completion and navigation.
@@ 5.2. Incomplete Metadata
We saw, consistent with prior work (Oreamuno et al., 2024; Jiang et al., 2023c, 2024c, d), that the documentation provided by model owners on Hugging Face is often incomplete. This is a trend not just for smaller, infrequently used projects but also for large projects with tens of thousands of downloads. For example, OpenAI’s clip-vit-large-patch14-336 model, with 5,827,027 downloads, has a model card that is almost entirely incomplete, save for basic hyper-parameter and framework version information (OpenAI, [n. d.]). Only 37.8% of models and 27.6% of datasets declared any kind of licensing information in a standardized, machine-readable way. We discuss the prevalence of declared components further in Section 6.
Some models may provide information on training data and architecture by linking to scientific papers that include these details, but such a practice introduces an additional hurdle in easily and programmatically obtaining that data, contributing to the ML reproducibility crisis (Jiang et al., 2024a). Ultimately, the reason why many models do not even mention the datasets used would require further investigation, including interviewing/surveying developers, which is out of the scope of this work.
@@
5.3. The “Unknown” license
The Hugging Face platform’s documentation provides a list of recognized licenses each with a unique, standardized short-hand identifier that can be utilized by users when creating their model/data cards (Face, 2024a). This list includes common Open Source Software (OSS), Creative Commons (CC), and ML-specific licenses as well as an “Other” catch-all category that encompasses less common, custom, or modified licenses not included in Hugging Face’s predefined license list. An “Unknown” licensing option is also provided, but there is no guidance provided as to when this option should be selected. Future work will need to explore when and why model owners select this option, particularly given that acknowledging that the license is “Unknown” suggests, at best, a lack of due diligence or understanding on the part of the owner or, at worst, that the model/dataset raises copyright infringement issues. We observed 4,419 (1.5%) models and 2,194 (4.5%) datasets that used this “Unknown” license tag, which inherently raises compliance challenges for dependent models.
5.4. Naming problems